2. Use the formula Pa(k) = PG-e(k) - PG/e(k) to find the chromatic polynomial of the following graph (take e as shown).

hurry up please
Answer #1

Sol: Here, first we draw the two graphs G-e and G/e. Then we find the chromatic polynomial of the two graphs. Finally, we apply the given formula to find the chromatic polynomial of the given graph G.Solh k -) k- G-e - Gle Here, k is the chromatic number, The number of possible colours for each of the vertices of the graphsin the chromatic polynomial of the given graph G من PG (k) = Pare (k) - Pale(K) =k? (k-1)} – k(k-1)} = (k-1)} (ki-k) = (k-1).Verifications k e k -1 k- k -) Pa(k) - K. (k-1). (k-1), (k-1)(k-1) -k(k-1)